# Coarse-graining nonequilibrium diffusions with Markov chains

### v2

@@ Abstract.
We investigate nonequilibrium steady-state dynamics in both continuous- and discrete-state stochastic processes. Our analysis focuses on planar diffusion dynamics and their coarse-grained approximations by discrete-state Markov chains. Using finite-volume approximations, we derive an approximate master equation directly from the underlying diffusion and show that this discretisation preserves key features of the nonequilibrium steady-state. In particular, we show that the entropy production rate of the approximation converges as the number of discrete states goes to the limit. These results are illustrated with analytically solvable diffusions and numerical experiments on nonlinear processes, demonstrating how this approach can be used to explore the dependence of entropy production rate on model parameters. Finally, we address the problem of inferring discrete-state Markov models from continuous stochastic trajectories. We show that discrete-state models significantly underestimate the true entropy production rate. However, we also show that they can provide tests to determine if a stationary planar diffusion is out of equilibrium. This property is illustrated with both simulated data and empirical trajectories from schooling fish.
@@ 1. Introduction
Many physical processes can be described by the interplay of deterministic forcing and random fluctuations. Such processes are typically described as diffusive and have been used to model a plethora of systems including Brownian motion [7], cell motility [41], climate dynamics [Franzke_O’Kane_2017], neural systems [29], and financial markets, amongst others. Mathematically, they can be formulated as Îto stochastic differential equations (SDE), taking the form
@@ 2.1. Diffusions in continuous-space: Stochastic differential equations and the Fokker-Planck equation
where $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a vector field that defines the deterministic forcing, known as the drift, and $D(x)=\frac{1}{2}\Sigma(x)\Sigma^{\top}(x):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times d}$ is the diffusion matrix, which is positive definite [58]. When using spatially-dependent diffusion, we will use the Îto convention for stochastic integration [58]. [^1] It is worth noting that the choice of convention for stochastic integration can lead to important differences in nonequilibrium properties like time-reversibility [76, 1]. The density of the process, $p(x,t)$, is defined as the probability density of the process attaining a value of $x\in\mathbb{R}^{d}$ at time $t>0$. Its dynamics are given by the Fokker-Planck (FP) equation
@@ 2.2. Discrete-state processes: Continuous-time Markov chains
When this condition is violated, the EPR of the NESS can be computed with the Schnakenberg formula [^3] At first glance, the discrete and continuous forms of the EPR may not appear to have much in common. However, both are defined as the relative entropy between the forward and backward path-space measures. This can be derived both measure-theoretically, and more heuristically by taking the derivative of the Shannon entropy of either the distribution, in the case of the CTMC, or of the entropy along a stochastic trajectory, in the case of the diffusion. For further details see Refs. [12, 34]. [32],
@@ 3.1. Non-Markovian approaches
Coarse-graining a Markovian diffusion with discrete-states leads to both memory effects and non-Markovian dynamics, in the same way that coarse-graining a Markov chain via state-aggregation induces memory in discrete-state dynamics [66, 14]. This typically leads to both non-exponential waiting times, as well as non-Markovian transitions [^4] It is important to note that these are separate phenomena. Whilst non-exponential waiting times imply a semi-Markov process, coarse-graining can also induce more disruptive memory effects where transition rates between states depend on the trajectory at large.. Previous approaches have used first-passage time analysis and asymptotic expansions in an attempt to obtain the waiting-time distributions of a discrete-state approximation of prototypical, univariate diffusions [23, 15, 25, 47]. Such an analysis is closely linked to approaches for estimating the EPR from semi-Markov processes [^5] A semi-Markov process is a discrete-state process with arbitrary waiting-time distributions [60]. using waiting-time distributions [44, 71]. Whilst such approaches more accurately describe the coarse-grained dynamics, they are typically limited in the complexity of the dynamics that can be applied to. Moreover, the statistical inference of semi-Markov processes from experimental data is more involved, hence data analytic methods typically use a Markovian model [2, 42, 53]. In discrete-time, it is simple to infer higher-order Markov models from sequence data, and thus reveal the effects of memory in real stochastic dynamics [61, 62].
@@ 3.2. A Markovian approximation
A numerical scheme yields a valid ME if $L_{(i,j),(k,l)}>0$ for all off-diagonal entries ($(k,l)\neq(i,j)$), and probability is conserved i.e. the columns of $L$ sum to 0. In Ref. [72], Strang considered the centred-difference approximation for the FP equation in 1D, which only yields a valid ME under step-size conditions. As a result, the problem of deriving a valid approximation is closely related to that of developing a structure-preserving scheme for the FP equation [31, 30]. We will adopt Scharfetter-Gummel discretisation [18], which alleviates step-size conditions.
@@ 3.3. The Scharfetter-Gummel discretisation
The Scharfetter-Gummel (SG) discretisation is a structure-preserving scheme for the FP equation that yields to the following approximation for the flux across the boundary,
@@
with analogous expressions for the remaining flux terms [18] (for a derivation, see App. A). This naturally leads to a ME equation of the form
@@
which are necessarily positive [^6] In cases, where the denominator vanishes, which occurs when the drift vanishes, then we can re-derive the SG discretisation to obtain $L_{(i,j),(i+1,j)}=D^{x}_{i+\frac{1}{2},j}/\Delta x^{2}$, and analogous terms for the other transitions [18].. Moreover, the diagonal rates are given by
@@
It is well-known that FVA conserve probability777i.e. the numerical solution at every time integrates to 1. This is equivalent to the derived Laplacian having columns that sum to 0. [39]. Moreover, the SG discretisation preserves the stationary distribution and many thermodynamic properties of the continuous solution [64], hence it is well-suited to this application.
@@ 4.1.2. The stochastic Hopf oscillator
The Hopf oscillator [^9] Also known as the Stuart-Landau oscillator [46]. (HO) is a nonlinear process given by the equations
@@ 4.1.3. Numerical experiments
We begin our numerical experiments by calculating the stationary distribution for the OU process defined in Eq. (45) for fixed values of $\theta=5$ and $\sigma=5$, whilst varying the discretisation size $\Delta x=\Delta y$ in a finite domain [^12] In this case, we use the domain $10,10]^{2}$ which we regard as sufficiently far from the centre of mass of the distribution that the no-flux boundaries do not significantly alter the behaviour of the discrete process. For a discussion of boundary conditions and their implementation see App. A.1.. In Fig. 4, we see that the stationary distribution of the coarse-grained dynamics preserves the symmetry of the underlying Gaussian distribution, even for larger step-sizes. As the step-size decreases, the distribution converges to the exact density [^13] To compare between a discrete distribution that must sum to 1, and a continuous density that must integrate to 1, we normalise the discrete distribution $\pi$ by the area of each volume $\Delta x^{2}$ i.e. the density in cell $i$ can be approximated discretely as $\pi_{i}/\Delta x^{2}$..
@@ 4.2.1. The stochastic van der Pol oscillator
where $B=I-\frac{\Delta t}{2}L$ and $A=I+\frac{\Delta t}{2}L$ for a time-step $\Delta t$ [39]. We begin with a initial Gaussian density with zero mean and isotropic deviation $\sigma=1$ and integrate up to $T=1.4$ with $\Delta t=0.2$ and $\Delta x=0.05$. As shown in Fig. 10, we see that the FP solution converges to the stationary distribution calculated directly.
@@
The shape and nature of the stationary distribution shifts with variations in parameters. In Fig. 11, we show three different parameter settings. Increasing $\mu$ increases the nonlinearity of the oscillation spreading the distribution over a larger limit cycle, whilst increased $\theta$ promotes circular rotation around the origin leading to a concentration of trajectories near the sinusoidal oscillation. Increasing $\sigma$ increases the diffusivity of trajectories and wider distribution.
Figure 11. NESS of the VDP oscillator. The shape and and nature of the stationary distribution shifts in response to changes in parameters. We illustrate this by calculating the stationary distribution of the VDP for an increased value of $\mu=3$, $\theta=5$ and $\sigma=2$ respectively. An increased value of $\mu$ promotes nonlinearity in the oscillation, whilst $\theta$ promotes circular rotation around the origin. Finally, $\sigma$ increases the diffusivity of the process, leading to a less peaked distribution. Additionally, we can investigate how the EPR varies with $\mu$ and $\theta$ in the coarse-grained dynamics. We find that $\theta$ increases the EPR whilst $\mu$ decreases it.
@@ 4.2.2. A pair of frustrated Kuramoto oscillators
The Kuramoto oscillator is a model that described the phase of interacting limit-cycle oscillators [37], and is typically used to model synchronisation. It has found a range of applications describing neural dynamics [54], the flashing of fireflies, or the voltage oscillations of Josephson junctions [73].
@@
which yields the constant stationary density $\pi(\theta)=1/2\bm{\pi}$, and stationary flux $J_{\text{ss}}(\theta)=\omega/2\bm{\pi}$ (Recall that we defined $\bm{\pi}$ to be ratio of a circle’s perimeter to its diameter, whilst $\pi$ is the stationary density). This leads to an EPR of $\Phi=2\omega^{2}/\sigma^{2}$, from Eq. (9), but constraining the integral to the domain $0,2\bm{\pi})$. This indicates that a single Kuramoto oscillator is in a NESS, with irreversibility driven by oscillation and mitigated by diffusion.
@@
where $\theta_{i}\in[0,2\bm{\pi})$ are the phases, $\omega_{i}>0$ are the natural frequencies, $\alpha$ is the frustration parameter, and $(A_{12},A_{21})$ are the coupling strengths. Fig. 12 shows the drift field for a range of different values. In the symmetric case, we see that the synchronised state is a stable equilibrium, which can be disrupted with asymmetry in the natural frequencies, interaction strengths or through frustration [^14] In fact, for $A_{12}=A_{21}$ and $\alpha=\pi/2$, the NESS is explicitly solvable. The stationary distribution is uniform, $\pi=1/4\bm{\pi}^{2}$ and the EPR is given by $\Phi=2(\omega_{1}^{2}+\omega_{2}^{2}+A_{12}^{2})/\sigma^{2}$. This shows that frustration can lead to a NESS, even in the case of symmetric coupling and no internal oscillations..
@@ 5. Statistical inference for discrete-state processes
where $N_{j\rightarrow k}$ is the number of observed transitions from state $j$ to state kk [^16] In practice, we must also avoid singularities, thus we employ a Bayesian prior. This means we add a ‘pseudo-count’ for all valid transitions [62].[6].
@@
Given $P$, the embedding problem is to find $L$ such that Eq. (63) is satisfied [11]. [^17] In our case, $P=\hat{P}$ is calculated from the observations using Eq. (62). Importantly, if $L$ solves the embedding problem for $\hat{P}$, then it is the MLE for the CTMC. However, it need not be unique as the matrix exponential is not injective. One simple condition that guarantees the existence and uniqueness of $L$ is
@@ 5.2. Determining if trajectories are from a NESS
Instead, in order to test if the dynamics of the stochastic data are in fact from a NESS, we analyse the trajectory as before, with one crucial difference. For each valid transition in the trajectory, we flip the direction of the original transition with probability 1/2, thus erasing the net flux in the system that results in entropy production and irreversibility. Then, the EPR of the genuine trajectory can be compared to a surrogate distribution obtained from different realisations of the surrogate model. Under the null hypothesis that the genuine trajectory is from an ESS, we can test if the trajectory is from a NESS using a one-sided $t$-test. We illustrate this approach on trajectories from the OU process with $\theta=5$, an irreversible process with $\Phi=25$, and $\theta=0$, a reversible process. Fig. 16 shows a comparison between the EPR of 100 trajectories sampled from the OU process compared with 100 surrogate models. At $\theta=5$, we find that the EPR of the original trajectories is significantly higher (one-sided independent $t$-test; ****, $p<0.0001$), than the surrogate trajectories, thus we can conclude that the original trajectories are sampled from a NESS. On the other hand, at $\theta=0$, we find that the original trajectories are not significantly higher than the surrogate models (one-sided independent $t$-test; ns, $p>0.05$), thus we must conclude, correctly, that the original process is reversible [^18] We note that in this case, the surrogate trajectories appear to have a significantly higher EPR than the original time-series. In theory, this shuffling approach should bring a NESS closer to detailed balance. On the other hand, if a trajectory is genuinely reversible, flipping a number of transitions is likely to increase the EPR. Although this effect should dissipate as the number of transitions increases..
@@ 5.2.1. An application to real-world trajectories: group-polarisation in schooling fish
Next, we investigate if the stationary dynamics of the schooling fish are best described as a NESS or an ESS. For each of the population sizes, we infer the EPR of the original trajectory as well as the EPR of 100 surrogate models, obtained using the shuffling procedure. Fig. 18 shows that, across population size, the EPR is not significantly higher for the original trajectories when compared to the surrogate model, from which we conclude that the process is not in a NESS. This result supports a well-known phenomenon in collective dynamics and active matter. Whilst collective dynamics are inherently active at a local level [74], as they dissipate energy in order for agents to adjust their states dynamically, their collective behaviour can be described by equilibrium steady-state dynamics respecting time-reversal symmetry [17] and maximum-entropy distributions [5, 51].

### v3

# arXiv:2511.05366v3[cond-mat.stat-mech] 02 Mar 2026

@@ Abstract.
We investigate nonequilibrium steady-state dynamics in both continuous- and discrete-state stochastic processes. Our analysis focuses on planar diffusion dynamics and their coarse-grained approximations by discrete-state Markov chains. Using finite-volume approximations, we derive an approximate master equation directly from the underlying diffusion and show that this discretisation preserves key features of the nonequilibrium steady-state. In particular, we show that the entropy production rate of the approximation converges as the number of discrete states goes to the limit. These results are illustrated with analytically solvable diffusions and numerical experiments on nonlinear processes, demonstrating how this approach can be used to explore the dependence of the entropy production rate on model parameters. Finally, we address the problem of inferring discrete-state Markov models from continuous stochastic trajectories. We show that discrete-state models significantly underestimate the true entropy production rate. However, we also show that they can provide tests to determine if a stationary planar diffusion is out of equilibrium. This property is illustrated with both simulated data and empirical trajectories from schooling fish.
@@ 1. Introduction
Many physical processes can be described by the interplay of deterministic forcing and random fluctuations. Such processes are typically described as diffusive and have been used to model a plethora of systems including Brownian motion [7], cell motility [41], climate dynamics [Franzke_O’Kane_2017], neural systems [29], and financial markets, amongst others. Mathematically, they can be formulated as Itô stochastic differential equations (SDE), taking the form
@@ 2.1. Diffusions in continuous-space: Stochastic differential equations and the Fokker-Planck equation
where $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a vector field that defines the deterministic forcing, known as the drift, and $D(x)=\frac{1}{2}\Sigma(x)\Sigma^{\top}(x):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times d}$ is the diffusion matrix, which is positive definite [58]. When using spatially-dependent diffusion, we will use the Itô convention for stochastic integration [58]. [^1] It is worth noting that the choice of convention for stochastic integration can lead to important differences in nonequilibrium properties like time-reversibility [76, 1]. The density of the process, $p(x,t)$, is defined as the probability density of the process attaining a value of $x\in\mathbb{R}^{d}$ at time $t>0$. Its dynamics are given by the Fokker-Planck (FP) equation
@@ 2.2. Discrete-state processes: Continuous-time Markov chains
When this condition is violated, the EPR of the NESS can be computed with the Schnakenberg formula [32], [^3] At first glance, the discrete and continuous forms of the EPR may not appear to have much in common. However, both are defined as the relative entropy between the forward and backward path-space measures. This can be derived both measure-theoretically, and more heuristically by taking the derivative of the Shannon entropy of either the distribution, in the case of the CTMC, or of the entropy along a stochastic trajectory, in the case of the diffusion. For further details see Refs. [12, 34].
@@ 3.1. Non-Markovian approaches
Coarse-graining a Markovian diffusion with discrete-states leads to both memory effects and non-Markovian dynamics, in the same way that coarse-graining a Markov chain via state-aggregation induces memory in discrete-state dynamics [66, 14]. This typically leads to both non-exponential waiting times, as well as non-Markovian transitions. [^4] It is important to note that these are separate phenomena. Whilst non-exponential waiting times imply a semi-Markov process, coarse-graining can also induce more disruptive memory effects where transition rates between states depend on the trajectory at large. Previous approaches have used first-passage time analysis and asymptotic expansions in an attempt to obtain the waiting-time distributions of a discrete-state approximation of prototypical, univariate diffusions [23, 15, 25, 47]. Such an analysis is closely linked to approaches for estimating the EPR from semi-Markov processes using waiting-time distributions [44, 71]. [^5] A semi-Markov process is a discrete-state process with arbitrary waiting-time distributions [60]. Whilst such approaches more accurately describe the coarse-grained dynamics, they are typically limited in the complexity of the dynamics that can be applied to. Moreover, the statistical inference of semi-Markov processes from experimental data is more involved, hence data analytic methods typically use a Markovian model [2, 42, 53]. In discrete-time, it is simple to infer higher-order Markov models from sequence data, and thus reveal the effects of memory in real stochastic dynamics [61, 62]
@@ 3.2. A Markovian approximation
A numerical scheme yields a valid ME if $L_{(i,j),(k,l)}>0$ for all off-diagonal entries ($(k,l)\neq(i,j)$), and probability is conserved i.e. the columns of $L$ sum to 0. In Ref. [72], Strang considered the centred-difference approximation for the FP equation in 1D, which only yields a valid ME under step-size conditions. As a result, the problem of deriving a valid approximation is closely related to that of developing a structure-preserving scheme for the FP equation [31, 30]. We will adopt the Scharfetter-Gummel discretisation [18], which alleviates step-size conditions.
@@ 3.3. The Scharfetter-Gummel discretisation
The Scharfetter-Gummel (SG) discretisation is a structure-preserving scheme for the FP equation that yields the following approximation for the flux across the boundary,
@@
with analogous expressions for the remaining flux terms [18] (for a derivation, see App. A). This naturally leads to a ME of the form
@@
which are necessarily positive. [^6] In cases, where the denominator vanishes, which occurs when the drift vanishes, then we can re-derive the SG discretisation to obtain $L_{(i,j),(i+1,j)}=D^{x}_{i+\frac{1}{2},j}/\Delta x^{2}$, and analogous terms for the other transitions [18]. Moreover, the diagonal rates are given by
@@
It is well-known that FVA conserve probability [39].777i.e. the numerical solution at every time integrates to 1. This is equivalent to the derived Laplacian having columns that sum to 0. Moreover, the SG discretisation preserves the stationary distribution and many thermodynamic properties of the continuous solution [64], hence it is well-suited to this application.
@@ 4.1.2. The stochastic Hopf oscillator
The Hopf oscillator (HO) is a nonlinear process given by the equations [^9] Also known as the Stuart-Landau oscillator [46].
@@ 4.1.3. Numerical experiments
We begin our numerical experiments by calculating the stationary distribution for the OU process defined in Eq. (45) for fixed values of $\theta=5$ and $\sigma=5$, whilst varying the discretisation size $\Delta x=\Delta y$ in a finite domain. [^12] In this case, we use the domain $10,10]^{2}$ which we regard as sufficiently far from the centre of mass of the distribution that the no-flux boundaries do not significantly alter the behaviour of the discrete process. For a discussion of boundary conditions and their implementation see App. A.1. In Fig. 4, we see that the stationary distribution of the coarse-grained dynamics preserves the symmetry of the underlying Gaussian distribution, even for larger step-sizes. As the step-size decreases, the distribution converges to the exact density. [^13] To compare between a discrete distribution that must sum to 1, and a continuous density that must integrate to 1, we normalise the discrete distribution $\pi$ by the area of each volume $\Delta x^{2}$ i.e. the density in cell $i$ can be approximated discretely as $\pi_{i}/\Delta x^{2}$.
@@ 4.2.1. The stochastic van der Pol oscillator
where $B=I-\frac{\Delta t}{2}L$ and $A=I+\frac{\Delta t}{2}L$ for a time-step $\Delta t$ [39]. We begin with an initial Gaussian density with zero mean and isotropic deviation $\sigma=1$ and integrate up to $T=1.4$ with $\Delta t=0.2$ and $\Delta x=0.05$. As shown in Fig. 10, we see that the FP solution converges to the stationary distribution calculated directly.
@@
The shape and nature of the stationary distribution shifts with variations in parameters. In Fig. 11, we show three different parameter settings. Increasing $\mu$ increases the nonlinearity of the oscillation spreading the distribution over a larger limit cycle, whilst increased $\theta$ promotes circular rotation around the origin leading to a concentration of trajectories near the sinusoidal oscillation. Increasing $\sigma$ increases the diffusivity of trajectories and leads to a wider distribution.
Figure 11. NESS of the VDP oscillator. The shape and nature of the stationary distribution shifts in response to changes in parameters. We illustrate this by calculating the stationary distribution of the VDP for an increased value of $\mu=3$, $\theta=5$ and $\sigma=2$ respectively. An increased value of $\mu$ promotes nonlinearity in the oscillation, whilst $\theta$ promotes circular rotation around the origin. Finally, $\sigma$ increases the diffusivity of the process, leading to a less peaked distribution. Additionally, we can investigate how the EPR varies with $\mu$ and $\theta$ in the coarse-grained dynamics. We find that $\theta$ increases the EPR whilst $\mu$ decreases it.
@@ 4.2.2. A pair of frustrated Kuramoto oscillators
The Kuramoto oscillator is a model that describes the phase of interacting limit-cycle oscillators [37], and is typically used to model synchronisation. It has found a range of applications describing neural dynamics [54], the flashing of fireflies, or the voltage oscillations of Josephson junctions [73].
@@
which yields the constant stationary density $\pi(\theta)=1/2\bm{\pi}$, and stationary flux $J_{\text{ss}}(\theta)=\omega/2\bm{\pi}$ (Recall that we defined $\bm{\pi}$ to be the ratio of a circle’s perimeter to its diameter, whilst $\pi$ is the stationary density). This leads to an EPR of $\Phi=2\omega^{2}/\sigma^{2}$, from Eq. (9), but constraining the integral to the domain $0,2\bm{\pi})$. This indicates that a single Kuramoto oscillator is in a NESS, with irreversibility driven by oscillation and mitigated by diffusion.
@@
where $\theta_{i}\in[0,2\bm{\pi})$ are the phases, $\omega_{i}>0$ are the natural frequencies, $\alpha$ is the frustration parameter, and $(A_{12},A_{21})$ are the coupling strengths. Fig. 12 shows the drift field for a range of different values. In the symmetric case, we see that the synchronised state is a stable equilibrium, which can be disrupted with asymmetry in the natural frequencies, interaction strengths or through frustration. [^14] In fact, for $A_{12}=A_{21}$ and $\alpha=\pi/2$, the NESS is explicitly solvable. The stationary distribution is uniform, $\pi=1/4\bm{\pi}^{2}$ and the EPR is given by $\Phi=2(\omega_{1}^{2}+\omega_{2}^{2}+A_{12}^{2})/\sigma^{2}$. This shows that frustration can lead to a NESS, even in the case of symmetric coupling and no internal oscillations.
@@ 5. Statistical inference for discrete-state processes
where $N_{j\rightarrow k}$ is the number of observed transitions from state $j$ to state kk[6]. [^16] In practice, we must also avoid singularities, thus we employ a Bayesian prior. This means we add a ‘pseudo-count’ for all valid transitions [62].
@@
Given $P$, the embedding problem is to find $L$ such that Eq. (63) is satisfied [11]. [^17] In our case, $P=\hat{P}$ is calculated from the observations using Eq. (62). Importantly, if $L$ solves the embedding problem for $\hat{P}$, then it is the MLE for the CTMC. However, it need not be unique as the matrix exponential is not injective. One simple condition that is sufficient for the existence and uniqueness of $L$ is
@@ 5.2. Determining if trajectories are from a NESS
Instead, in order to test if the dynamics of the stochastic data are in fact from a NESS, we analyse the trajectory as before, with one crucial difference. For each valid transition in the trajectory, we flip the direction of the original transition with probability 1/2, thus erasing the net flux in the system that results in entropy production and irreversibility. Then, the EPR of the genuine trajectory can be compared to a surrogate distribution obtained from different realisations of the surrogate model. Under the null hypothesis that the genuine trajectory is from an ESS, we can test if the trajectory is from a NESS using a one-sided $t$-test. We illustrate this approach on trajectories from the OU process with $\theta=5$, an irreversible process with $\Phi=25$, and $\theta=0$, a reversible process. Fig. 16 shows a comparison between the EPR of 100 trajectories sampled from the OU process compared with 100 surrogate models. At $\theta=5$, we find that the EPR of the original trajectories is significantly higher (one-sided independent $t$-test; ****, $p<0.0001$), than the surrogate trajectories, thus we can conclude that the original trajectories are sampled from a NESS. On the other hand, at $\theta=0$, we find that the original trajectories are not significantly higher than the surrogate models (one-sided independent $t$-test; ns, $p>0.05$), thus we must conclude, correctly, that the original process is reversible. [^18] We note that in this case, the surrogate trajectories appear to have a significantly higher EPR than the original time-series. In theory, this shuffling approach should bring a NESS closer to detailed balance. On the other hand, if a trajectory is genuinely reversible, flipping a number of transitions is likely to increase the EPR. Although this effect should dissipate as the number of transitions increases.
@@ 5.2.1. An application to real-world trajectories: group-polarisation in schooling fish
Next, we investigate if the stationary dynamics of the schooling fish are best described as a NESS or an ESS. For each of the population sizes, we infer the EPR of the original trajectory as well as the EPR of 100 surrogate models, obtained using the shuffling procedure. Fig. 18 shows that, across population size, the EPR is not significantly higher for the original trajectories when compared to the surrogate model, from which we conclude that the process is not in a NESS. This result supports a well-known phenomenon in collective dynamics and active matter. Whilst collective dynamics are inherently active at a local level [74], as they dissipate energy in order for agents to adjust their states dynamically, their collective behaviour can be described by equilibrium dynamics respecting time-reversal symmetry [17] and maximum-entropy distributions [5, 51].